OpenStudy (anonymous):

whats the next number to the pattern . -0.6,-3,-15,-75?

OpenStudy (charlotte123):

-0.6 * 5 = 3 -3 * 5 = -15 -15 * 5 = -75 So - To find the next number to the pattern - You multiply -75 by 5. -75 * 5 = -375 ^ The Answer